40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 25 16:05:51 2022, Yehonathan, Update, SUS, Trying to fix some SUS PR2_Step_Response_Test_2022-08-25_16-23.pdfPR3_Step_Response_Test_2022-08-25_18-37.pdf
    Reply  Tue Sep 6 17:39:40 2022, Paco, Update, SUS, LO1 LO2 AS1 AS4 damping loop step responses LO1_Step_Response_Test_2022-09-06_17-19.pdfLO2_Step_Response_Test_2022-09-06_17-30.pdfAS1_Step_Response_Test_2022-09-06_17-53.pdfAS4_Step_Response_Test_2022-09-06_18-16.pdf
       Reply  Tue Nov 8 10:23:42 2022, Anchal, Update, SUS, MC3 damping loop step responses MC3_Step_Response_Test_2022-11-08_09-48.pdfMC3_Step_Response_Test_2022-11-08_10-19.pdf
Message ID: 17133     Entry time: Tue Sep 6 17:39:40 2022     In reply to: 17105     Reply to this: 17241
Author: Paco 
Type: Update 
Category: SUS 
Subject: LO1 LO2 AS1 AS4 damping loop step responses 

I tuned the local damping gains for LO1, LO2, AS1, and AS4 by looking at step responses in the DOF basis (i.e. POS, PIT, YAW, and SIDE). The procedure was:

  1. Grab an ndscope with the error point signals in the DOF basis, e.g. C1:SUS-LO1_SUSPOS_IN1_DQ
  2. Apply an offset to the relevant DOF using the alignment slider offset (or coil offset for the SIDE DOF) while being careful not to trip the watchdog. The nominal offsets found for this tuning are summarized below:
Alignment/Coil Step sizes
  POS PIT YAW SIDE
LO1 800 300 300 10000
LO2 800 300 400 -10000
AS1 800 500 500 20000
AS4 800 400 400 -10000
  1. Tune the damping gains until the DOF shows a residual Q with ~ 5 or more oscillations.
  2. The new damping gains are below for all optics and their DOFs, and Attachments #1-4 summarize the tuned step responses as well as the other DOFs (cross-coupled).
Local damping gains
  POS PIT YAW SIDE
LO1 10.000 5.000 3.000 40.000
LO2 10.000 3.000 3.000 50.000
AS1 14.000 2.500 3.000 85.000
AS4 15.000 3.100 3.000 41.000

Note that during this test, FM5 has been populated for all these optics with a BounceRoll (notches at 16.6, 23.7 Hz) filter, apart from the Cheby (HF rolloff) and the 0.0:30 filters.

Attachment 1: LO1_Step_Response_Test_2022-09-06_17-19.pdf  79 kB  Uploaded Tue Sep 6 18:59:22 2022  | Hide | Hide all
LO1_Step_Response_Test_2022-09-06_17-19.pdf LO1_Step_Response_Test_2022-09-06_17-19.pdf LO1_Step_Response_Test_2022-09-06_17-19.pdf LO1_Step_Response_Test_2022-09-06_17-19.pdf
Attachment 2: LO2_Step_Response_Test_2022-09-06_17-30.pdf  76 kB  Uploaded Tue Sep 6 19:01:09 2022  | Hide | Hide all
LO2_Step_Response_Test_2022-09-06_17-30.pdf LO2_Step_Response_Test_2022-09-06_17-30.pdf LO2_Step_Response_Test_2022-09-06_17-30.pdf LO2_Step_Response_Test_2022-09-06_17-30.pdf
Attachment 3: AS1_Step_Response_Test_2022-09-06_17-53.pdf  77 kB  Uploaded Tue Sep 6 19:02:46 2022  | Hide | Hide all
AS1_Step_Response_Test_2022-09-06_17-53.pdf AS1_Step_Response_Test_2022-09-06_17-53.pdf AS1_Step_Response_Test_2022-09-06_17-53.pdf AS1_Step_Response_Test_2022-09-06_17-53.pdf
Attachment 4: AS4_Step_Response_Test_2022-09-06_18-16.pdf  77 kB  Uploaded Tue Sep 6 19:17:13 2022  | Hide | Hide all
AS4_Step_Response_Test_2022-09-06_18-16.pdf AS4_Step_Response_Test_2022-09-06_18-16.pdf AS4_Step_Response_Test_2022-09-06_18-16.pdf AS4_Step_Response_Test_2022-09-06_18-16.pdf
ELOG V3.1.3-